M 340L 54185 Fall ’98 T Th 11:00-12:30 ESB 333

Instructor: C. Friedman RLM 11.140 471-5161 office hrs: T Th 2:30-3:30pm or by appt. (Web page: http://www.ma.utexas.edu/˜ friedman)

Text and Syllabus: Linear Algebra and Its Applications, 2nd ed. , David C. Lay, Addison Wesley, 1997 Syllabus: Selected sections of Chapters I-VI.

Prerequisites: M408D

Exams: We will have exams on 9/29 (class 10; Tues.), 11/3 (class 20; Tues.), 12/3 (Class 28; Thurs.) (No “Final exam”). All exam grades will be curved.

* Exam grades will be scaled (if the unscaled average is ≤ 75) to some value in the 75 to 80 range. MAKEUPS WILL BE GIVEN VERY RELUCTANTLY

Homework: Homework problems will be assigned each Tuesday in class and will be due the next Tuesday. A serious effort should be made on these exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, all homework together will be counted as a grade equivalent to one exam grade. (Homework assignments will be listed on my web page - see above.)

Final Grades: In computing a final grade, I will count the three tests and one total homework grade equally. (So there will be four grades to average.)